OpenStudy (anonymous):

Find the x-intercept and y-intercept of each line. -y=12-4x

OpenStudy (anonymous):

x intercept is achieved by replacing y by 0 y intercept is achieved by replacing x by 0 can you do it now ?

OpenStudy (anonymous):

No... N

OpenStudy (anonymous):

*No...not really

OpenStudy (anonymous):

Okay,Wait. Is this right? -y=12-4x -y=12-4(0) Then divide both sides by -y and the y-intercept is -12?

OpenStudy (anonymous):

yes .. y intercept is -12

OpenStudy (anonymous):

Okay..But I don't know how to find the x-intercept.

OpenStudy (anonymous):

replace y by 0

OpenStudy (anonymous):

So..It's 3?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Thanks a ton!

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What should I do to 2x+y=7 ?

OpenStudy (anonymous):

same thing

OpenStudy (anonymous):

replace y by 0 for x intercept replace x by 0 for y intercept